Ballistics experts are testing projectiles. Projectile A has mass m and is fired with a velocity 2v. Projectile B has mass 2m and is fired with velocity v. Projectile C has mass m and is fired at a speed v. Which projectile has the greatest kinetic energy?

(A) Projectile A

(B) Projectile B

(C) Projectile C

(D) Both A and B have equal kinetic energy




1
Expert's answer
2021-05-18T19:43:52-0400

"K.E =1\/2mv^2"

for A, "K.E=1\/2m(2v)^2=2mv^2"

for B, "K.E=1/2\times2m\times v^2=mv^2"

for C, "K.E=1\/2mv^2=0.5mv^2"

thus A has the greatest kinetic energy